### v3

# Buyer-Initiated Auction Mechanism for Data Redemption in Machine Unlearning

Abstract
The rapid growth of artificial intelligence (AI) has raised privacy concerns over user data, leading to regulations like the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA). With the essential toolbox provided by machine unlearning, AI service providers are now able to remove user data from their trained models as well as the training datasets, so as to comply with such regulations. However, extensive data redemption can be costly and degrade model accuracy. To balance the cost of unlearning and the privacy protection, we propose a buyer-initiated auction mechanism for data redemption, enabling the service provider to purchase data from willing users with appropriate compensation. This approach does not require the server to have any a priori knowledge about the users’ privacy preference, and provides an efficient solution for maximizing the social welfare in the investigated problem.
Index Terms:
Data privacy, data redemption, machine unlearning, auction mechanism.
I Introduction
The rapid development of artificial intelligence (AI) and machine learning (ML) technologies over the past decade has brought about a prospery of intelligent applications and a new era of digital economy, recently supplemented by the explosive blowout of large language models (LLMs). As the feul of this new techno-economic paradigm, high-quality data has become a valuable asset and in thirst demand. In practice, the data used for training ML models is often collected from the users of the applications. However, the data collection process may raise privacy concerns, as the data may contain sensitive information about the users, and the models therewith trained may leak or abuse such information against the users’ will.
Being aware of these privacy concerns, data regulations have been established in various countries and regions, represented by the General Data Protection Regulation (GDPR) in the European Union [1], the California Consumer Privacy Act (CCPA) in California [2], and the Personal Information Protection Law (PIPL) in China [3]. Such regulations generally require the service providers to obtain the users’ consent before collecting their data, and to provide the users with the right to delete their data afterwards.
To enable removing user’s data from the server in both senses of the training dataset and the trained model, the technologies of machine unlearning have been developed and attracted much attention [4]. Despite the technical success, however, an excessive unlearning of user data can be not only computationally costly for the AI service provider, but also significantly degrading the model’s accuracy, which also harm the users’ quality of experience [5]. Therefore, a rigid unlearning policy may be the long run lead against both interests of the service provider and the users.
To address this issue, a new data redemption framework in on call for, which shall balance the cost of unlearning and the privacy protection. One of the appealing approaches is data monetization, i.e., the AI service provider encourages its users not to redeem their data by offering them a compensation. This allows the users to flexibly choose between reserving and selling their data upon their own will and interest, while the service provider benefits from the reduced cost of unlearning and better model accuracy.
The crucial challenge in this framework is to determine the price of data. Although data pricing has been extensively studied in the context of data marketplaces [6], existing works generally deal with the scenario of active data trading for model training, focusing on the quality of data and the model performance, while the data redemption and unlearning scenario is barely explored. This substential difference leads to an untolarable inapplicability of the existing pricing models in the data redemption context. For example, some works such like [7] simply ignores the privacy loss of users when they sell data to the server, while other works such like [8], though taking privacy into account, generally do not capture the computational cost caused by unlearning.
A piorneering work on the topic of pricing in data redemption is presented in [9], where the authors propose a novel incentive mechanism for machine unlearning. Their model considers the server’s cost of unlearning regarding both computation and accuracy degradation, and also takes into account the users’ privacy utility. A two-stage optimization problem is formulated, where on the first stage the server determines an optimal unit price for user data, and on the second stage every user determin its own optimal amount of data to sell.
Despite the novelty and inspiration of the work in [9], there are still some limitations in it. First, the server’s optimal pricing relies on certain assumptions in the distribution of users’ privacy parameters, which may not hold in practical scenarios. Second, the two-stage pricing mechanism always returns one uniform price for every data batch for all users. Considering the heterogeneity of users’ privacy concerns and the non-linearity of user’s privacy utility regarding the redeemed data amount, this is possibly leading to a suboptimum.
In this work, we propose to invoke buyer-initiated ascending auction mechanism to address the above limitations. In a multi-round auction, the server gradually raises the unit price for purchasing data, and each user independently decides how much data it is willing to sell at the offered price. After the users claiming, the server determines how much data it is willing to buy at the current price, before announcing the new price for the next round or closing the auction. This auction mechanism enables a flexible and progressive pricing to convey the non-linear cost/utility functions of both the server and the users, and requires no prior knowledge on the server side about the distribution of users’ privacy parameters.
The remainding part of this paper is organized as follows. In Section II, we present the system model, including the cost and utility functions of the server and the users, respectively. In Section III, we analyze the server’s incentive to purchase data and the users’ incentive to sell data in an arbitrary auction round. In Section IV, we propose the buyer-initiated auction mechanism, and discuss the optimal decision strategies for the server and the users. Subsequently, we numerically evaluate our proposed approach with simulation campaigns in Sec. V, before finally concluding the paper in Section VI.
II System Model
II-A Server Cost Model
Considering the cost model in [9] that $C(x)=\alpha A(x)+\beta T(x)$, where $x=\sum\limits_{i=i}^{I}x_{i}$, $x_{i}$ is the amount of data redemption for user $i\in\mathcal{I}$, $A(x)$ is the accuracy degradation of the model caused by unlearning, and $T(x)$ the computing time for executing the unlearning task. Note that $\alpha,A,\beta,T$ are all non-negative. Defining the total data amount for each user $i$ as $d_{i}$, and their sum $d=\sum\limits_{i=1}^{I}d_{i}$, according to [9]:
$A(x)$ $=A_{1}a^{A_{2}x}-A_{3},$ (1)
$T(x)$ =$=\begin{cases}0&\text{if }x=0,\\ T_{0}(d-x)&\text{if }x\in(0,d].\end{cases}$ (2)
From the server’s perspective it is more convenient to focus on the amount of data to keep, i.e. to buy from the users, rather than the amount of data to redeem. So we define that $y_{i}=d_{i}-x_{i}$ and $y=d-x$, and rewrite the cost function as
$C(y)$ $=\alpha A(y)+\beta T(y),$ (3)
$A(y)$ $=A_{1}e^{A_{2}(d-y)}-A_{3},$ (4)
$T(y)$ =$=\begin{cases}T_{0}y&\text{if }y\in[0,d),\\ 0&\text{if }y=d.\end{cases}$ (5)
II-B User Privacy Model
Following [9] we consider that for each user $i\in\mathcal{I}$ the utility of privacy is captured by
$U_{i}(x_{i})=\lambda_{i}\ln(x_{i}+1),$ (6)
where $\lambda_{i}$ is its privacy parameter. Similarly, focusing on the amount of data to keep on the server, we rewrite it As
$U_{i}(y_{i})=\lambda_{i}\ln(d_{i}-y_{i}+1),$ (7)
which is concave and monotonically decreasing w.r.t. $y_{i}$.
III Incentive Analyses
III-A Server’s Incentive to Purchase Data
The server has an incentive to keep as much data as to minimize the cost of unlearning. We denote this target amount of data to keep as $y_{\mathrm{max}}$. Note that $C(y)$ has one and only one jump discontinuity at $y=d$, for simplification of discussion we first focus on the continuous interval $y\in[0,d)$, and leave the case of $y=d$ for later discussion in Sec. IV-D. Thus, $C(y)$ is in general a convex function over $y\in[0,d)$, while its monoticity, depending on the specific parameter settings, may have three diffferent cases, as shown in Fig. 1:
  1. 1):
    monotonically increasing,
  2. 2):
    monotonically decreasing, or
  3. 3):
    non-monotone.
Figure 1: Different cases of the server’s cost function.
In the first case, trivially, the server will not keep any data and $y_{\mathrm{max}}=0$. In the second case, the server tend to keep all data and $y_{\mathrm{max}}\to d^{-}$. In the third case, $y_{\mathrm{max}}$ can be obtained by looking for the zero derivative:
$\left.\frac{\text{d}C(y)}{\text{d}y}\right|_{y_{\mathrm{max}}}$ $=(-\alpha A_{1}A_{2}\ln a)a^{A_{2}(d-y_{\mathrm{max}})}+\beta T_{0}=0,$ (8)
$\Rightarrow y_{\mathrm{max}}$ $=d+\frac{1}{A_{2}}\log_{a}\left(\frac{\alpha A_{1}A_{2}\ln a}{\beta T_{0}}\right).$ (9)
Now assume that the server has already managed to keep $y<y_{\mathrm{max}}$ amount of data, and offers to buy more data from users at a price $B$ per unit amount of data. The server’s demand $\eta$ will be a function of the condition $(y,B)$, which is:
$\eta(y,B)=\underset{\delta\in[0,y_{\mathrm{max}}-y]}{\arg\max}\left[C(y)-C(y+\delta)-B\delta\right]$ (10)
$=$ $\underset{\delta\in[0,y_{\mathrm{max}}-y]}{\arg\max}\left[\alpha A_{1}e^{A_{2}(d-y)}(1-e^{-A_{2}\delta)})-(\beta T_{0}+B)\delta\right].$
Especially, noticing that $C(y)$ is always monotonic over the interval $y\in[0,y_{\mathrm{max}})$, there is a unique maximum price $B_{\mathrm{all}}(y)$ to support a “buy all” strategy:
$\begin{split}&B_{\mathrm{all}}(y)=\frac{C(y_{\mathrm{max}})-C(y)}{y_{\mathrm{max}}-y}\\ =&\frac{\alpha A_{1}e^{A_{2}d}\left(e^{-A_{2}y_{\mathrm{max}}}-e^{-A_{2}y}\right)}{y_{\mathrm{max}}-y}+\beta T_{0}.\end{split}$ (11)
With $y$ data already bought from the users, given any price $B\leqslant B_{\mathrm{all}}(y)$, the server will buy as much data from the users as possible, until it reaches the maximum amount $y_{\mathrm{max}}$ of data to keep. Note that in case of $y$-increasing cost $C$ (Case 1 in Fig. 1), $B_{\mathrm{all}}\leqslant 0$ and $y_{\mathrm{max}}=0$.
III-B User’s Incentive to Sell Data
Given the price $B$ offered by the server, each user $i$ can decide how much data it will sell to the server, based on how much data $y_{i}$ it has already sold previously. The change of privacy utility of selling $\delta_{i}$ more data is
$\begin{split}&\Delta U_{i}(y_{i},\delta_{i})=U_{i}(y_{i}+\delta_{i})-U_{i}(y_{i})\\ =&\lambda_{i}\ln\left(1-\frac{\delta_{i}}{d_{i}-y_{i}+1}\right),\end{split}$ (12)
so the minimum price to convince user $i$ to sell $\delta_{i}$ more data must compensate this utility loss:
$\begin{split}&B_{\text{min},i}(y_{i},\delta_{i})=\frac{-\Delta U_{i}(y_{i},\delta_{i})}{\delta_{i}}\\ =&\frac{\lambda_{i}}{\delta_{i}}\ln\left(1+\frac{\delta_{i}}{d_{i}-y_{i}+1-\delta_{i}}\right).\end{split}$ (13)
It is trivial that for all $y_{i}\in[0,d_{i})$ and $\delta_{i}\geqslant 0$:
$\frac{\partial}{\partial y_{i}}B_{\text{min},i}(y_{i},\delta_{i})$ $>0,$ (14)
$\frac{\partial}{\partial\delta_{i}}B_{\text{min},i}(y_{i},\delta_{i})$ $>0,$ (15)
so there is a minimal price for user $i$ to sell any amount of data in addition to the already-sold $y_{i}$:
$\lim\limits_{\delta_{i}\to 0}B_{\text{min},i}(y_{i},\delta_{i})=\frac{\lambda_{i}}{d_{i}-y_{i}+1},$ (16)
and a minimal price for user $i$ to sell all the remaining data:
$B_{\text{all},i}(y_{i})=B_{\text{min},i}(y_{i},d_{i}-y_{i})=\frac{\lambda_{i}}{d_{i}-y_{i}}\ln(d_{i}-y_{i}+1).$ (17)
Moreover, given that user $i$ has already sold $y_{i}$ in priori, there is a maximal amount of data $q_{i}$ it is willing to sell to the server at price $p$, which we call $i$’s supply:
$\begin{split}&q_{i}(y_{i},B)=\underset{\delta\in[0,d_{i}-y_{i}]}{\arg\max}\left[p\delta+\Delta U_{i}(y_{i},\delta)\right].\\ =&\underset{\delta\in[0,d_{i}-y_{i}]}{\arg\max}\left[p\delta+\lambda_{i}\ln\left(1-\frac{\delta}{d_{i}-y_{i}+1}\right)\right].\end{split}$ (18)
IV Multi-Round Auction Mechanism
IV-A Buyer-Initiated English Auction
To balance the server’s cost of unlearning and the users’ privacy utility, we propose a buyer-initiated ascending auction mechanism, which can be briefly summarized as follows:
  1. 1.
    At the beginning of each round $t$, the server updates its price $B^{t}$ for buying unit data from the users. We focus in this study on the ascending auction format, where the data price is increasing w.r.t. time, i.e., for all $t<\tau$, $B^{t}<B^{\tau}$. It also estimates its demand for data $\eta^{t}=\eta(y^{t},B^{t})$, which is unknown to the users. If the demand drops to zero, the auction is terminated. Otherwise, the price is announced to all users.
  2. 2.
    Each user then independently decides how much data it can supply at this price.
  3. 3.
    The server purchases the data from all users who are willing to sell at the announced price. If the total supply exceeds its demand, it will only purchase its demanded amount of data. Otherwise, it will buy all data supplied.
  4. 4.
    If the server has collected the target amount of data $y_{\mathrm{max}}$, the auction is terminated. Otherwise, $t\leftarrow t+1$, and go back to step 1) for the next round.
For the sake of low communication overhead and privacy protection, our approach is to let the users make decisions independently from each other. Thus, each user $i$ is considered blind in the sense that at each period $t$, $i$ only knows the current data price $B^{t}$, its own privacy parameter $\lambda_{i}$, its own data amount $d_{i}$, and its own history of data selling $(y_{i}^{\tau})_{\tau<t}$. As we are considering an ascending auction, the price announced by the server along the time assembles an ascending series $(B^{\tau})_{1\leqslant\tau<\infty}$.
IV-B Users’ Selling Strategy
Noticing the $t$-monotonically increasing price $B^{t}$, when a user has incentive to sell data in round $t$, it has also the alternative option of keeping the data to sell it at higher price in a future round. Herewith we discuss the optimal selling strategy of users in such multi-round auction.
Suppose the auction is now at the beginning of round $t$. Each user $i$’s current remaining data that can be sold is $x_{i}^{t}\geqslant 0$. Given the information that $i$ knows, $i$ forms a belief $\mu_{i}^{t}\in[0,1]$, which represents a subjective probability that the auction is terminated after round $t$. Thus, $\mu_{i}^{t}$ is a function about server’s remaining demand $\eta^{t}$ at $t$, and the supply $q_{j}^{t}$ of all users $j\in\mathcal{I}$.
Moreover, at $t$, since $i$ only knows its own history $(y_{i}^{\tau})_{\tau<t}$, and that $y_{j}^{\tau}$ is non-decreasing about $\tau$ for all $j\in\mathcal{I}$, $\mu_{i}^{t}$ is increasing with $i$’s own sold data $y_{i}^{t}$, and the collection $(\mu_{i}^{\tau})_{1\leqslant\tau<\infty}$ is (point-wisely) decreasing with $\tau$, for any $(y_{i}^{\tau})_{1\leqslant\tau<\infty}$.
Let $t_{i}$ be the first period that $i$ decides to sell, where according to the data price $B^{t_{i}}$, the optimal amount of selling for $i$ is automatically given by (18). Denote this amount of data be $q_{i}^{t_{i}}=q_{i}(0,B^{t_{i}})>0$. If $q_{i}^{t_{i}}=d_{i}$, it is straightfoward. Otherwise, supposing $q_{i}^{t_{i}}<d_{i}$, by the section of $t_{i}$ we have
$\mu_{i}^{t_{i}}(q_{i}^{t_{i}})\cdot V(q_{i}^{t_{i}},B^{t_{i}})\geqslant[1-\mu_{i}^{t_{i}}(0)]\cdot V(q_{i}^{t_{i}+1},B^{t_{i}+1}),$ (19)
where $V(y,B)$ is $i$’s benefit by selling the amount of data $y$ with price $B$, according to his utility. Thus, Eq. (19) implies that the expected payoff of selling at $t_{i}$ is higher than the the expected payoff of keeping data at $t_{i}$. In other words, as long as the users are blind about the server’s demand and other users’ trade record, they always perform the greedy strategy in selling data.
Furthermore, rearranging Eq. (19), we have
$\frac{\mu_{i}^{t_{i}}(q_{i}^{t_{i}})}{[1-\mu_{i}^{t_{i}}(0)]}\geqslant\frac{V(q_{i}^{t_{i}+1},B^{t_{i}+1})}{V(q_{i}^{t_{i}},B^{t_{i}})}.$ (20)
With the aforementioned monotonicity of $\mu_{i}$, we also have
$\frac{\mu_{i}^{t_{i}+1}(q_{i}^{t_{i}+1})}{[1-\mu_{i}^{t_{i}+1}(0)]}\geqslant\frac{\mu_{i}^{t_{i}}(q_{i}^{t_{i}})}{[1-\mu_{i}^{t_{i}}(0)]}.$ (21)
By the definitions of $q$ and $V$, $i$’s marginal benefit is decreasing, and hence
$\frac{V(q_{i}^{t_{i}+1},B^{t_{i}+1})}{V(q_{i}^{t_{i}},B^{t_{i}})}\geqslant\frac{V(q_{i}^{t_{i}+2},B^{t_{i}+2})}{V(q_{i}^{t_{i}+1},B^{t_{i}+1})}.$ (22)
Thus, we can obtain that
$\frac{\mu_{i}^{t_{i}}(q_{i}^{t_{i}})}{[1-\mu_{i}^{t_{i}}(0)]}\geqslant\frac{V(q_{i}^{t_{i}+2},B^{t_{i}+2})}{V(q_{i}^{t_{i}+1},B^{t_{i}+1})}.$ (23)
Eq. (23) implies, if $i$ starts to sell data at $t_{i}$, it will also choose to sell data in all future rounds, until having all its data sold.
IV-C Handling the Oversupply
Due to the wide-sense $B$-monotonicity of $\eta$ (decreasing) and $q_{i}$ (increasing) for all $i\in\mathcal{I}$, when the announced price $B^{t}$ is sufficiently high, it may occur the case of $\eta^{t}<\sum\limits_{i=1}^{I}q_{i}^{t}$, i.e. the total supply from users exceeds the server’s demand. In this case, the server shall allocate its demand $\eta$ to the different selling users.
While neither the unit price $B^{t}$ nor the demand $\eta^{t}$ is influenced by the allocation, the server’s purchase is independent therefrom, either. However, since the different users are of different characteristics $(y_{i},\lambda_{i})$, the total payoff of users will be depending on the specific allocation of purchase. To perform an optimal allocation that maximizes the users’ payoff, the server needs full knolwdge about the privacy parameter $\lambda_{i}$ of each selling user $i$, which is assumed unavailable in our system model. Therefore, here we propose four privacy-knowledge-free strategies to handle such oversupply issue:
  1. 1.
    Major sellers first: The server will purchase data from users in a greedy manner and descending order upon their current supply $q_{i}^{t}$, until the demand is satisfied.
  2. 2.
    Minor sellers first: The server will purchase data from users in a greedy manner and ascending order upon their current supply $q_{i}^{t}$, until the demand is satisfied.
  3. 3.
    Proportional: The server will purchase data from all users, allocating its demand $\eta^{t}$ to every single selling user proportionally to their current supply $q_{i}^{t}$.
  4. 4.
    Random order: The server will purchase data from users in a greedy manner and random order.
Remark that the oversupply will certainly trigger the satisfaction of server’s optimal amount of data to keep $y_{\mathrm{max}}$, so it will only occur once, i.e., in the last round of the auction.
IV-D Handling the Discontinuity in Cost
It must be recalled that the server’s cost function $C(y)$ has a jump discontinuity at $y=d$, which is so far not considered in the above analyses.
When $y_{\mathrm{max}}<d$, due to this jump discountinuity, the actual server’s cost of unlearning all user data of amount $d$ is $C(d)=\alpha A(d)$, which can be significantly lower than $C(y_{\mathrm{max}})$. Thus, after achieving $y^{t}=y_{\mathrm{max}}$, the server may still have an incentive to see if it worth buying all the remaining data from the users. Nevertheless, since the server does not have full knowledge about every user’s privacy parameter, it cannot simply estimate the optimal price to offer. It is neither rational to simply continue purchasing more data at prices higher than the marginal cost of unlearning. To address this issue, we propose a post-auction procedure, where after buying $y_{\mathrm{max}}$ amount of data, the server will still keep announcing new prices to the users like if the auction continues, but not purchasing. Instead, it observes if all users are willing to sell all remaining data at the new price. Only if so, the server will purchase all remaining data at the last announced price, and terminates the post-auction step. Otherwise, no purchase will be made even if the supply is non-zero, and the server simply updates the price. This procedure continues until the price $B^{t}$ is raised so high that the cost to buy all remaining data exceeds the unlearning cost saved therewith, i.e. when $B^{t}\cdot\left(d-y_{\mathrm{max}}\right)>C\left(y_{\mathrm{max}}\right)-C(d)$.
IV-E Proposed Algorithm
Summarizing the above discussions, we propose the algorithm for the buyer-initiated ascending auction in Algorithm 1. Especially, for practical implementation in real use scanrios, we consider discrete user data sets with unit size of $\Delta d$, so that all trades of user data can only be executed by amount of integer times of $\Delta d$.
1 Input: $a,A_{1},A_{2},A_{3},T_{0},\alpha,\beta,B^{0},\Delta B,[d_{1},d_{2},\dots,d_{I}],\Delta d$,
2 Initialize: $y^{0}\leftarrow 0$, $t\leftarrow 0$, $y_{\mathrm{max}}$ w.r.t. Eq. (9), $\eta^{0}$ w.r.t. Eq. (10)
3 while $\eta^{t}>0$ do // Auction phase
4       Update $q_{i}^{t}$ for all $i\in\mathcal{I}$ w.r.t. Eq. (18)
5       $q_{i}^{t}\leftarrow\left\lfloor q_{i}^{t}/\Delta d\right\rfloor*\Delta d$
6       if $\sum\limits_{i=1}^{I}q_{i}^{t}<\eta^{t}$ then
7             $y_{i}^{t+1}\leftarrow y_{i}^{t}+q_{i}^{t}$ for all $i\in\mathcal{I}$
8
9       else
10             Update $y_{i}^{t+1}$ for all $i\in\mathcal{I}$ upon the oversupply-handling strategy
11
12       end if
13      $B^{t+1}\leftarrow B^{t}+\Delta B$
14       $\eta^{t+1}\leftarrow\eta\left(y^{t+1},B^{t+1}\right)$
15       $t\leftarrow t+1$
16
17 end while
18while $B^{t}\cdot\left(d-y^{t}\right)\leqslant C\left(y^{t}\right)-C(d)$ do // Post-auction phase
19       Update $q_{i}^{t}$ for all $i\in\mathcal{I}$ w.r.t. Eq. (18)
20       if $\sum\limits_{i=1}^{I}q_{i}^{t}=d-y$ then
21             $y_{i}^{t+1}\leftarrow d_{i}$ for all $i\in\mathcal{I}$
22
23       end if
24      $B^{t+1}\leftarrow B^{t}+\Delta B$
25       $t\leftarrow t+1$
26
27 end while
Algorithm 1 Buyer-Initiated English Auction
V Numerical Evaluation
V-A Simulation Setup
To verify our proposed approach, we conducted numerical simulation campaigns. The simulation setup is shown in Table I, following the specifications of [9].
TABLE I: Simulation Setup
Parameter Value Remark
$I$ $10$ Number of users
$d_{i},\forall i\in\mathcal{I}$ 6000 Data amount per user
$\lambda_{i},\forall i\in\mathcal{I}$ $\sim\mathcal{U}(0.5,29.5)$ User privacy
$\Delta d$ $1$ Unit data amount
$B^{0}$ $0.001$ Initial price
$\Delta B$ $0.001$ Price step
System $N$ $1000$ Monte Carlo runs
$[a,A_{1},A_{2},A_{3}]$ $[e,0.1,3.33\times 10^{-5},0]$ Accuracy parameters
$T_{0}$ $2.85\times 10^{-4}$ Time factor
Cost $[\alpha,\beta]$ [1500,1] Weight factors
V-B Oversupply Handling Strategies
To compare the performance of the four oversupply-handling strategies proposed in Sec. IV-C, we carried out Monte Carlo tests. For each run, we measured the the server’s payoff:
$\xi_{\mathrm{s}}=C^{T}-C(0)-\sum\limits_{t=0}^{T}B^{t},$ (24)
where $T$ is the index of round when the procedure terminates, the users’ total payoff
$\xi_{\mathrm{u}}=\sum\limits_{i=1}^{I}\left[\sum\limits_{t=1}^{T}B^{t-1}(y_{i}^{t}-y_{i}^{t-1})+U_{i}\left(y_{i}^{T}\right)-U_{i}(0)\right],$ (25)
and the social welfare
$\xi=\xi_{\mathrm{s}}+\xi_{\mathrm{u}}.$ (26)
The simulation results are shown in Table II. As we have expected, the selection of oversupply-handling strategy does not affect the server’s payoff at all (difference within $0.2\%$), while the minor-seller-first strategy outperforms the other three regarding the users’ payoff and social welfare by a very slight superiority (within $2\%$). This is not suprising as the oversupply handling is triggered only in the last auction round, which takes only a small portion of the total auction procedure.
TABLE II: Comparing the oversupply-handling strategies
Major-first Minor-first Prop. Rand.
Server’s payoff 684.8 683.6 684.1 684.9
Users’ payoff 1133.5 1158.2 1150.2 1134.0
Social welfare 1818.4 1841.8 1834.3 1818.9
V-C Baseline Methods
To comparatively benchmark the performance of our proposed auction approach against state-of-the-art solutions, we also implemented three baseline methods:
  1. 1.
    Do-not-redeem (DNR): This is a naive baseline where none of the users redeems any data, i.e., the server can keep all the user data for free, and therefore does not need to unlearn or purchase at all. However, all users undergo maximum privacy loss.
  2. 2.
    GDPR: This is the rigid privacy solution, where no data trade option is provided, and the server must unconditionally unlearn all the user data upon request.
  3. 3.
    Bi-stage pricing (BSP): This is the solution proposed in [9], where the server solves an optimal universal price and announce it to all users, each user then independently decides how much data to sell at that price.
It is also worth to introduce the concept of informed ratio following [9], which describes the ratio of users that are aware of their rights to ask for data redemption. The uninformed users will be dealed with the DNR solution, no matter which solution is used for the informed users.
V-D Benchmark Test
We conducted a benchmark test to compare the performance of our proposed auction approach with the aforementioned baseline methods. Specifically, the auction approach was configured to handle oversupply with the minor-sellers-first strategy. We tested all methods at different informed ratios from $0\%$ to $100\%$ with $10\%$ step length, and measured the server’s payoff, users’ payoff, and social welfare. The results of Monte Carlo tests are shown in Fig. 2.
Figure 2: Performance of different data redemption solutions.
As we can see, our proposed auction approach outperforms all the other methods in terms of social welfare. While the social welfare brought by the rigid GDPR solution drops at high informed ratio due to the high unlearning cost, both our auction approach and the BSP solution exhibit an social welfare steadily increasing along with the informed ratio. Particularly, in comparison to BSP, our method significantly improves the server’s payoff at the cost of a less reduction in users’ payoff, leading to its superiority in the overall social welfare. This shall be explained in two aspects:
  1. 1.
    The multi-round auction mechanism allows the server to explore the users’ non-linear privacy function over multiple intervals of data trading amount, which provides a closer approximation of the privacy utility loss by the payment. Compared to the BSP solution that provides only one optimum average price, the auction mechanism is more cost-profitable for the buyer (server), but less profitable for the users.
  2. 2.
    The convexity of of the server’s cost function $C$ about the amount of data under keep $y$ guanrantees that the gain of the server’s payoff is larger than the loss of the users’ payoff.
VI Conclusions
In this paper, we have proposed a buyer-initiated ascending auction framework for data redemption in machine unlearning. We have analyzed the incentives of buying and selling data of the server and users, respectively, and derived the optimal selling strategy of users. Compared with state-of-the-art baselines, our proposed solution does not require the server to possess any knowledge about the users’ privacy parameter, and yet achieving a higher social welfare, which is demonstrated by numerical simulations.
Acknowledgment
This work is supported by the German Federal Ministry of Education & Research in the project Open6GHub (16KISK003K / 16KISK004). B. Han (bin.han@rptu.de) is the corresponding author.